OpenStudy (anonymous):

What are the root(s) of the function (x) = -4x + 8? x= 2 x= 0, 3 x= -4 x= 3

OpenStudy (anonymous):

Putnam?

OpenStudy (anonymous):

put f(x)=0; you will get -4x+8=0 x=8/4=2 answer A

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

irshadno1 are u human lol